This Question: 1 pt 8 of 25 (6 complete) This Test: 25 pts possi The following information is from the 2019 records of Fast Lane Racing Gear Accounts receivable, December 31, 2019 Allowance for Bad Debts, December 31, 2019 prior to adjustment Net credit sales for 2019 Accounts written off as uncollectible during 2019 Cash sales during 2019 $45,000 (debit) 1,700 (debit) 178,000 16,000 28,000 Bad debts expense is estimated by the aging- of-receivables method. Management estimates that $7,000 of accounts receivable will be uncollectible. Calculate the ending balance of Allowance for Bad Debts, after the adjustment for bad debts expense, at December 31, 2019 ○ A. $10,400 O B. $5,340 O C. $6,000 O D. $7,000 Click to select your answer
Answer #1

Answer:

Option (D) 7,000 is correct answer

Ending Balance Allowance for Bad Debt = Bad Debt Expenses - Debit Balance of Allowance for Bad Debt
Note: Bad Debt Expenses = 7,000+1,700 = 8,700
Ending Balance of Allowance for Bad Debt = 8,700 -1,700
Ending Balance of Allowance for Bad Debt = 7,000